Operating wavelength : 439. nm
Absorbance of solution 1 - Standard Dichromate-H2O : 0.589 a.u.
Absorbance of solution 2 - Standard Dichromate-Ethanol (Standard) : 0.437 a.u.
Absorbance of solution 3 - Standard Dichromate-Ethanol (Defendant) : 0.381 a.u.
Mass of dichromate in Solution 1 : 6.25 mg


II. Calculated data:


Enter the following calculated data from the calibration of the Breath-Analyzer:


Mass of dichromate in Solution 2 :
Answer #1

Potassium dichromate produces a yellow solution when dissolved in acidic water solution. An acidic solution of dichromate reacts with ethanol, the Cr(VI) species is reduced to a Cr(III) species, which gives green tint to aqueous solution.

The overall reaction is as follows:

3 CH3CH2OH + Cr2O72– + 8 H+ ===> 3 CH3CHO + 2 Cr3+ + 7 H2O

The decrease in absorbance is associated with the oxidation of ethanol.

From Beer's law

Absorbance, A = ε *c*b = 0.589

we replace c by mass of dichromate :

ε *b = 0.589 / 6.25 mg = 0.09424 mg-1

Absorbance of solution 2 : Standard Dichromate-Ethanol (Standard)  = 0.437

thus mass of dichromate in Solution 2 : 0.437 / 0.09424 mg-1 = 4.64 mg